A car is travelling at a constant speed of 25 m/s. It begins to accelerate at a rate of 2.7 m/s2 for 7.1 seconds.
(a) How far does it travel during this time?
1
Expert's answer
2020-11-17T13:03:51-0500

The final speed of the car is


"v_f=v_i+at=25+2.7\cdot7.2=44\text{ m/s}."